# Leveraging temporal features of the divergence measure to detect chaos in conservative systems

### v1

@@ Abstract.
The recurrence-based divergence quantifier ($DIV$), traditionally applied to dissipative systems, is shown here to be an effective finite-time chaos indicator for conservative dynamics. We benchmark its performances against the well-established fast Lyapunov indicator (FLI), focusing on the standard map, a canonical model of Hamiltonian chaos. Through extensive numerical simulations on moderately long orbits, we find strong agreement between $DIV$ and FLI, supporting the reported correlation between the divergence of recurrences and positive Lyapunov exponents. Additionally, our study sheds more light into asymptotic time properties of $DIV$ by revealing distinct power laws for regular and chaotic trajectories, both in the original and reconstructed phase spaces. In particular, for regular dynamics, $DIV$ decays in average with the time $N$ as $1/N$, mirroring the decay rate of the maximal Lyapunov exponent. For chaotic trajectories, $DIV$ decreases at a much slower rate in average, close to $1/\sqrt{N}$. This scaling insight opens new avenues for characterizing chaos from time series. Our numerical results thus demonstrate $DIV$ to be a computationally viable and theoretically rich tool for chaos detection in conservative systems.
@@ 1. Introduction
The story of recurrence plot (RP) and the beginning of its methodological developments is often traced back to the seminal paper of Eckmann et al., (1987). Let be $\gamma$ a finite orbit, $\gamma=\{z_{0},z_{1},z_{2},\cdots,z_{n}\}$ of points in $\mathbb{R}^{d}$, $d\geq 1$, of some dynamical systems. The $\{z_{i}\}_{i}$ might be the results of mapping iterations, flow discretisation or observed states (measurements). The 010-1 recurrence matrix $R_{\gamma}=(r_{i,j})_{0\leq i,j\leq n}$ is constructed from the recurrence coefficient
(1) $r_{i,j}=\Theta\big(\left\lVert z_{i}-z_{j}\right\rVert-\epsilon\big),$
where $\Theta$ is the Heaviside function, $\epsilon$ is a real cutoff parameter, and $\left\lVert\bullet\right\rVert$ is a norm in $\mathbb{R}^{d}$. When $r_{i,j}=1$, $z_{i}$ and $z_{j}$ are recurrence points, which means that they are $\epsilon$ close. The matrix $R_{\gamma}$ always contains the line of identity as trivial recurrence points. A RP, as indicated by the name, is in its root a graphical tool. It consists of visualising the binary matrix $R_{\gamma}$ with recurrence points encoded as dark pixels. Interestingly enough, dynamical properties of the orbit (such as its oscillatory nature, the presence of drifting components, the presence of extreme or rare events, etc.) are encoded in different structures and textures of the RP (Eckmann et al., (1987); Webber Jr and Zbilut, (1994)). (Several typologies obtained on the standard map, the model on which we base our analysis, will be exemplified later in Sect. 2.) In addition to qualitative graphical assessments of RP, which can be difficult to interpret or intuit visually, several recurrence variables have been introduced by Webber Jr and Zbilut, (1994) and developed in greater depth over the years (Marwan et al.,, 2007), a field termed recurrence quantification analysis (RQA). Such variables include the recurrence rate $RR$ (percentage of recurrence points), the determinism $DET$ (proportion of recurring points forming diagonal lines) or lengths of vertical lines, to name but a few. In this contribution, we focus specifically on the RQA variable related to the length of the longest diagonal lines $\ell_{\max}$ in a RP (leaving aside the line of identity and its vicinity), and more precisely its inverse — another RQA variable — the divergence $DIV$ (Webber Jr and Zbilut,, 1994), defined as
@@
It has long been reported $\ell_{\max}$ (or $DIV$) to be in direct proportion to the largest Lyapunov exponent (Eckmann et al.,, 1987; Zbilut and Webber,, 1992; Webber Jr and Zbilut,, 1994; Trulla et al.,, 1996), as heuristically understood here. Assume the trajectory to be embedded into some $\mathbb{R}^{d}$ and that $r_{i,j}$ is a recurrence point. It follows that $z_{j}$ is $\epsilon$ close to $z_{i}$. If we interpret $\delta_{i,j}=z_{i}-z_{j}$ as a deviation vector and follow its time evolution, in case of chaotic dynamics, we expect the quantity $\{\left\lVert\delta_{i+k,j+k}\right\rVert\}_{k}$ to grow exponentially fast in average with $k$ at a rate dictated by the largest Lyapunov exponent. Thus, the probability of $z_{i+k,j+k}$ still being an $\epsilon$ neighbour of $z_{i,j}$ (a recurrent point) decreases with time $k$ very fast. In other words, the length of the diagonal that stems from the observed times $(i,j)$ is expected to be small. A more rigorous link between diagonal lines and chaos is made apparent when looking at the distribution of diagonal lines of length $\ell$ and the second-order Rényi entropy (also called correlation entropy), $K_{2}$ (Grassberger and Procaccia,, 1983). For purely deterministic systems $K_{2}=0$ and for stochastic systems $K_{2}\to+\infty$. Deterministic chaos is characterised by finite values, $K_{2}>0$. It turns out that $K_{2}$ can be estimated directly from RPs as presented in Thiel et al., (2003). An algorithm to allow its numerical estimation is detailled by Asghari et al., (2004) and Marwan et al., (2007). The procedure involves essentially three major steps. Firstly, one need to compute the cumulative distribution of the diagonal lines for several choice of $\epsilon$. Secondly, one need to identify automatically a scaling region and clusters with $\epsilon$. Finally, a last step involving a linear regression leads to the estimation of $K_{2}$. Although some contributions have applied the RP and RQA frameworks to conservative dynamics (see in particular Zou et al., 2007a ; Zou et al., 2007b ; Zou et al., (2016) or Asghari et al., (2004); Kovács, (2019, 2020) in the context of gravitational $n$-bodies like dynamics), their developments have been mostly driven by dissipative dynamics. This might be partly due to the important application of RPs and RQAs to the field of time series analysis (see Marwan, (2023) for a bibliographical view of RPs and RQAs), the raise of chaos in low-dimensional dynamical systems, and the desire to analyse complex systems from the nonlinear time series perspective (intimately connected with phase space reconstruction and delay embedding theorem, see, e.g.,Abarbanel et al., (1993)). This contribution has two main objectives: first, to further apply the methodology and potential of RPs to conservative problems by focusing on the standard map model, paradigm of Hamiltonian chaos; and second, to propose a methodology for using the specific $DIV$ metric as an indicator of chaos. Indeed, our work provides numerical evidences that the more heuristic divergence $DIV$ furnishes a simple and robust workaround of the more cumbersome automatisation of the estimation of $K_{2}$, and might be reliably used to distinguish between regular and chaotic motions [^1] The landscape of chaos indicators and complexity measures developed over the last decades is rich, and still active. Among the various existing methods, it is rather customary to distinguish between variational methods, relying on the Jacobian associated to the dynamics (or, an estimation of it), and orbit based diagnostics. The divergence $DIV$ has the advantage to belong to the later family; yet, its complexity is $\mathcal{O}(N^{2})$, $N$ being the length of the orbit. A non-exhaustive list of indicators include the Lyapunov exponent (Benettin et al.,, 1976), the fast Lyapunov indicator and variations of it (Froeschlé et al.,, 1997; Fouchard et al.,, 2002), the mean-exponential growth of nearby orbits (MEGNO, Cincotta and Simó, (2000)), the smaller alignement index and its generalisation (SALI and GALI respectively, see Skokos, (2001); Skokos et al., (2007)), frequency analysis (Laskar,, 1993), topological braid chaos methods (Thiffeault,, 2005), the 010-1 test (Gottwald and Melbourne,, 2004), Lagrangian descriptors methods based on lengths of orbits (Daquin et al.,, 2022; Hillebrand et al.,, 2022) (requiring the trajectory and the knowledge of nearby trajectories), Birkhoff averages (Sander and Meiss,, 2020). .
@@ 2.1. The standard map
The standard map is a 2-dimensional area preserving map paradigmatic of Hamiltonian chaos. The map is obtained as the Poincaré map of the periodically kicked rotator model (Chirikov,, 1979). Given a point $(p_{0},\theta_{0})$ on the two dimensional torus $\mathbb{T}=[0,2\pi]^{2}$, the dynamics for $n\geq 0$ reads
@@
where $K$ is the nonlinearity parameter. When $K=0$, the dynamics is easily described. The “action” $p$ is constant and the angle $\theta$ evolves linearly with time. When $K\neq 0$, resonances grow in size and chaos manifests (Meiss,, 2008).
@@ 2.2. Computation of the divergence DIVDIV
The divergences $DIVs$ follow from the maximal line length in the diagonal direction of the RPs. The RPs and the RQAs measures are computed using the Julia language and specific packages (Bezanson et al.,, 2017; Datseris,, 2018). Several parameters are involved during the computation of the RPs matrices, provided here for the sake of reproducibility. Unless otherwise stated, the norm used in Eq. (1) is the Euclidean norm. To minimise arbitrariness, the parameter $\epsilon$ follows from a normalised value, a fixed recurrence rate (RR),
@@
here set to $RR=5\%$. For RPs computed up to time $N$, the longest diagonal line is the line of identity (trivial diagonal of recurrences). This line is excluded from the RPs, together with its vicinity (we set the Theiler window to 2). The smallest diagonal line considered in RPs is $\ell_{\min}=2$. The fig. 1 is an illustrative composite plot showing the phase space of the standard map at $K=1$ together with several RPs for four orbits with distinct properties: an orbit immersed within the main resonant island, an orbit in a secondary resonance, an orbit experiencing stickiness and finally one orbit experiencing large scale chaos.
Figure 1. (Left) The phase portrait of the standard map at $K=1$ for $N=500$ iterations. (Right) Recurrence plots associated to four representative trajectories shown in color in the phase space; red - quasi-periodic orbit $(\theta_{0},p_{0})=(3.5,1.0),$ blue - orbit trapped in a secondary resonance $(\theta_{0},p_{0})=(3.85,1.7),$ green - large scale chaotic orbit $(\theta_{0},p_{0})=(3.3,1.8),$ black - sticky orbit $(\theta_{0},p_{0})=(3.14,2.215).$ The different textures in the RPs serve as a basis for recurrence quantification analysis. In particular, we focus here on the divergence metric DIVDIV, associated to the non-trivial longest diagonal line in a given RP.
@@ 2.3. The fast Lyapunov indicator
A strictly positive largest finite time Lyapunov exponent (FTLE) is usually considered as signature of deterministic chaos (see Skokos, (2009) for a review), i.e.,nearby trajectories diverge exponentially fast in average. Here, to assess the presence of chaos, we instead compute a closely related quantity, namely the fast Lyapunov indicator, hereafter denoted FLI (Froeschlé et al.,, 1997). The FLI is a well established variational chaos indicator, valid for discrete and continuous systems. Contrarily to the FTLE, the FLI does not average the growths of the tangent vector over time, an average that impedes its fast convergence. Given a smooth mapping $x_{n+1}=M(x_{n})$, $n\in\mathbb{N}$, let us denote by $v_{0}$ a unitary deviation vector. The tangent map dynamics reads
@@
There are several definitions of the FLI literature, and in the following we follow the one of Lega and Froeschlé, (2001) that suppresses oscillations,
@@
For regular orbits, FLI grows as $\mathcal{O}(\log N)$ with time $N$. For chaotic orbits, the norm of the deviation vector grows exponentially fast, and thus the FLI tends to grow as $\mathcal{O}(N)$. Those two distinct time evolutions allow the detection of chaos in a short time (Guzzo and Lega,, 2023).
@@ 3. Time behavior of DIVDIV
Aiming to harness the potential of the $DIV$ measure as chaos indicator, we first need to ensure that the metric reacts differently on regular and chaotic trajectories. We have investigated time asymptotic properties of $DIV$ on three scenarios, detailed subsequently, by probing parametrically the dynamics of the standard map for many different initial conditions. Our parametric study focuses on properties in the original $2$ dimensional phase space, but also in reconstructed phase spaces when considering generated time series. Quite interestingly, our numerical results highlight clear distinct power laws underpinning the time behavior of the spatial average $\langle DIV\rangle$ for regular or chaotic trajectories. In the following, the final time $N$ is set to $N=25,000$, considered as here as an “asymptotic time.”

### v2

# Leveraging temporal features of the divergence quantifier of recurrence plot to detect chaos in conservative systems

@@ Abstract.
The recurrence-based divergence quantifier ($DIV$), traditionally applied to dissipative systems, is shown here to be an effective finite-time chaos indicator for conservative dynamics. We benchmark its performances against the well-established fast Lyapunov indicator (FLI), focusing on the standard map, a canonical model of Hamiltonian chaos. Through extensive numerical simulations on moderately long orbits, we find strong agreement between $DIV$ and FLI, supporting the reported correlation between the divergence of recurrences and positive Lyapunov exponents. Additionally, our study sheds more light into asymptotic time properties of $DIV$ by revealing distinct power laws on regular and chaotic components, both in the original and reconstructed phase spaces. In particular, on a regular component, the space average of $DIV$ decays with the time $N$ as $1/N$, mirroring the decay rate of the maximal Lyapunov exponent. On chaotic components, the space average of $DIV$ decreases at a much slower rate, close to $1/\sqrt{N}$. This scaling insight opens new avenues for characterizing chaos from time series. Our numerical results thus demonstrate $DIV$ to be a computationally viable and theoretically rich tool for chaos detection in conservative systems.
@@ 1. Introduction
The story of recurrence plot (RP) and the beginning of its methodological developments is often traced back to the seminal paper of jpEc87. Let be $\gamma$ a finite orbit of points in $\mathbb{R}^{d}$, $d\geq 1$, $\gamma=\{z_{0},z_{1},z_{2},\cdots,z_{n}\}$, of some dynamical systems. The $\{z_{i}\}_{i}$ might be the results of mapping iterations, flow discretisation or observed states (measurements). The 010-1 recurrence matrix $R_{\gamma}=(r_{i,j})_{0\leq i,j\leq n}$ is constructed from the recurrence scalar
(1) $r_{i,j}=\Theta\big({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon_{i}}-\left\lVert z_{i}-z_{j}\right\rVert\big),$
where $\Theta$ is the Heaviside function, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon_{i}$} is a real cutoff parameter, and $\left\lVert\bullet\right\rVert$ is a norm in $\mathbb{R}^{d}$. When $r_{i,j}=1$, $z_{i}$ and $z_{j}$ are recurrence points, which means that they are $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon_{i}$} close. The matrix $R_{\gamma}$ always contains the line of identity as trivial recurrence points and tends to be farily symmetrical [^1] When the cutoff parameter $\epsilon$ is independent of the time indices $i$, the recurrence plot is symmetric with respect to the first diagonal.. A RP, as indicated by the name, is in its root a graphical tool. It consists of visualising the binary matrix $R_{\gamma}$ with recurrence points encoded as dark pixels. Interestingly enough, dynamical properties of the orbit (such as its oscillatory nature, the presence of drifting components, the presence of extreme or rare events, etc.) are encoded in different structures and textures of the RP (jpEc87; chWe94). (Several typologies obtained on the standard map, the model on which we base our analysis, will be exemplified later in Sec. 2.) In addition to qualitative graphical assessments of RP, which can be difficult to interpret or intuit visually, several recurrence variables have been introduced by chWe94 and developed in greater depth over the years (nMa07-PR), a field termed recurrence quantification analysis (RQA). Such variables include the recurrence rate $RR$ (percentage of recurrence points), the determinism $DET$ (proportion of recurring points forming diagonal lines) or lengths of vertical lines, to name but a few.
In this contribution, we focus specifically on the RQA variable related to the length of the longest diagonal lines $\ell_{\max}$ in a RP (leaving aside the line of identity and its vicinity), and more precisely its inverse — another RQA variable — the divergence $DIV$ (chWe94), defined as
@@
It has long been reported $\ell_{\max}$ (or $DIV$) to be in direct proportion to the largest Lyapunov exponent (jpEc87; jpZb92; chWe94; lTr96), as heuristically understood here. Assume the trajectory to be embedded into some $\mathbb{R}^{d}$ and that $r_{i,j}$ is a recurrence point. It follows that $z_{i}$ is $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon_{i}$} close to $z_{j}$. If we interpret $\delta_{i,j}=z_{i}-z_{j}$ as a deviation vector and follow its time evolution, in case of chaotic dynamics, we expect the quantities $\{\left\lVert\delta_{i+k,j+k}\right\rVert\}_{k}$ to grow exponentially fast in average with $k$ at a rate dictated by the largest Lyapunov exponent. Thus, the probability of $z_{i+k,j+k}$ to be still an $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon_{i}$} neighbour of $z_{i,j}$ (a recurrent point) decreases with time $k$ very fast. In other words, the length of the diagonal that stems from $(i,j)$ is expected to be small.
A more rigorous link between diagonal lines and chaos is made apparent when looking at the distribution of diagonal lines of length $\ell$ and the second-order Rényi entropy (also called correlation entropy), $K_{2}$ (pGr83). For purely deterministic systems $K_{2}=0$ and for stochastic systems $K_{2}\to+\infty$. Deterministic chaos is characterised by finite values, $K_{2}>0$. It turns out that $K_{2}$ can be estimated directly from RPs as presented in mTh03. An algorithm to allow its numerical estimation is detailled by nAs04 and nMa07-PR. The procedure involves essentially three major steps. Firstly, one need to compute the cumulative distribution of the diagonal lines for several choice of $\epsilon$. Secondly, one need to identify automatically a scaling region and clusters with $\epsilon$. Finally, a last step involving a linear regression leads to the estimation of $K_{2}$.
Although several contributions have applied the RP framework and RQA quantifiers to conservative dynamics (see, for examples, yZo07; yZo07-1; yZo16; mPa22; mrSa23, or nAs04; tKo19; tKo20 in the context of gravitational $n$-bodies like dynamics), their developments have been mostly driven by dissipate dynamics. This might be partly due to the important application of RPs and RQAs to the field of time series analysis (see nMa23 for a bibliographical view of RPs and RQAs), the raise of chaos in low-dimensional dynamical systems, and the desire to analyse complex systems from the nonlinear time series perspective (intimately connected with phase space reconstruction and delay embedding theorem, see, e.g.,hAb93). This contribution has two main objectives: first, to further apply the methodology and potential of RPs to conservative problems by focusing on the standard map model, paradigm of Hamiltonian chaos; and second, to propose a methodology for using the specific $DIV$ quantifier as an indicator of chaos. Indeed, our work provides numerical evidences that the more heuristic divergence $DIV$ furnishes a simple and robust workaround of the more cumbersome automatisation of the estimation of $K_{2}$, and might be reliably used to distinguish between regular and chaotic motions.
The landscape of chaos indicators and complexity measures developed over the last decades is rich, and still an active field of research. Among the various existing methods, it is rather customary to distinguish between variational methods, relying on the Jacobian associated to the dynamics (or, an estimation of it), and orbit based diagnostics. The divergence $DIV$ has the advantage to belong to the later family; yet, its complexity is $\mathcal{O}(N^{2})$, $N$ being the length of the orbit. A non-exhaustive list of indicators include the Lyapunov exponent (gBe76), the fast Lyapunov indicator and variations of it (cFr97; mF022), the mean-exponential growth of nearby orbits (MEGNO, pCi00), the smaller alignement index and its generalisation (SALI and GALI respectively, see hSk01; hSk07), frequency analysis (jLa93), topological braid chaos methods (jlTh05), the 010-1 test (gGo04), Lagrangian descriptors methods based on lengths of orbits (jDa22; mHi22; jDa23) (requiring the trajectory and the knowledge of nearby trajectories), Birkhoff averages (eSa20).
@@ 2.1. The standard map
The standard map is a two-dimensional area preserving map paradigmatic of Hamiltonian chaos. The map is obtained as the Poincaré map of the periodically kicked rotator model (bCh79). Given a point $(p_{0},\theta_{0})$ on the two dimensional torus $\mathbb{T}=[0,2\pi]^{2}$, the dynamics for $n\geq 0$ reads
@@
where $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K>0$} is the nonlinearity parameter. When $K=0$, the dynamics is easily described. The “action” $p$ is constant and the angle $\theta$ evolves linearly with time. When $K>0$, resonances grow in size and chaos manifests (jMe08).
@@ 2.2. Computation of the divergence DIVDIV
The divergences $DIVs$ follow from the maximal line length in the diagonal direction of the RPs. The RPs and the RQAs measures are computed using the Julia language and specific packages (Julia-2017; Julia-2018). Several parameters are involved during the computation of the RPs matrices, provided here for the sake of reproducibility. Unless otherwise stated, the norm used in Eq. (1) is the Euclidean norm. To minimise arbitrariness, the parameters $\epsilon_{i,j}$ follows from a normalised value, a fixed recurrence rate (RR),
@@
here set to $RR=5\%$. For RPs computed up to time $N$, the longest diagonal line is the line of identity (trivial diagonal of recurrences). This line is excluded from the RPs, together with a Theiler window of size $w=2$, i.e.,a band of size $w=2$ around it. Mathematically, this reads
(6) $R_{i,j}=0,\,\,i,j=1,\cdots,N,$
whenever $|i-j|\leq w.$ The smallest diagonal line $\ell_{\min}$ considered in RPs is $\ell_{\min}=2$. The Fig. 1 is an illustrative composite plot showing the phase space of the standard map at $K=1$ together with several RPs for four orbits with distinct properties: an orbit immersed within the main resonant island, an orbit in a secondary resonance, an orbit experiencing stickiness and finally one orbit experiencing large scale chaos.
Figure 1. (Left) The phase portrait of the standard map at $K=1$ for $N=500$ iterations. (Right) Recurrence plots associated to four representative trajectories shown in color in the phase space; red - quasi-periodic orbit $(\theta_{0},p_{0})=(3.5,1.0),$ blue - orbit trapped in a secondary resonance $(\theta_{0},p_{0})=(3.85,1.7),$ green - large scale chaotic orbit $(\theta_{0},p_{0})=(3.3,1.8),$ black - sticky orbit $(\theta_{0},p_{0})=(3.14,2.215).$ The different textures in the RPs serve as a basis for recurrence quantification analysis. In particular, we focus here on the divergence quantifier DIVDIV, associated to the non-trivial longest diagonal line in a given RP.
@@ 2.3. The fast Lyapunov indicator
A strictly positive largest finite time Lyapunov exponent (FTLE) is usually considered as signature of deterministic chaos (see hSk09 for a review), i.e.,nearby trajectories diverge exponentially fast in average. Here, to assess the presence of chaos, we instead compute a closely related quantity, namely the fast Lyapunov indicator, hereafter denoted FLI (cFr97). The FLI is a well established variational chaos indicator, valid for discrete and continuous systems. Contrarily to the FTLE, the FLI does not average the growths of the tangent vector over time, an average that impedes its fast convergence. Given a smooth mapping $x_{n+1}=M(x_{n})$, $n\in\mathbb{N}$, let us denote by $v_{0}$ a unitary deviation vector. The tangent map dynamics reads
@@
There are several definitions of the FLI literature, and in the following we follow the one of eLe01 that suppresses oscillations,
@@
For regular orbits, FLI grows as $\mathcal{O}(\log N)$. For chaotic orbits, the norm of the deviation vector grows exponentially fast, and thus the FLI evolves linearly with time as $\mathcal{O}(N)$. Those two distinct time evolution allow the detection of chaos in a short time (mGu23). The time evolution of the FLI is exemplified for regular and chaotic orbits of the standard map in Fig. 2.
Figure 2. Time evolution of the fast Lyapunov indicator for a chaotic, circulational and librational orbit (respectively in green, blue, red) of the standard map with $K=0.6$. The distinct $\mathcal{O}(N)$ and $\mathcal{O}(\log N)$ growth rates allow the separation of chaotic and regular orbits in a short time.
@@ 3. Time behavior of DIVDIV
Aiming to harness the potential of the $DIV$ measure as chaos indicator, we first need to ensure that the quantifier reacts differently on regular and chaotic trajectories. We have investigated time asymptotic properties of $DIV$ on three scenarios, detailed subsequently, by probing parametrically the dynamics of the standard map for many different initial conditions. Our parametric study focuses on properties in the original two-dimensional phase space, but also in reconstructed phase spaces when considering generated time series. Quite interestingly, our numerical results highlight clear distinct power laws underpinning the time behavior of the spatial average $\langle DIV\rangle$ for regular or chaotic trajectories. In the following, the final time $N$ is set to $N=25,000$, considered as here as an “asymptotic time.”